1.
Rewrite the following quadratic equation in standard form: y = (x + 3)2 - 11
A. y = x - 2
B. y = 7x - 2
C. y = x2 + 6x + 20
D. y = x2 + 6x - 2

Answer by Cromlix(4381)

y = (x + 3)2 - 11
y = x^2 + 6x + 9 - 11
y = x^2 + 6x - 2 (D)
Hope this helps.
:-)